Calculate Marginal Utility and Average Utility from the information given in the below


table:


Quantity Consumed Total Utility


1 20


2 35


3 47


4 55


5 60

1
Expert's answer
2021-11-19T09:03:01-0500

Marginal utility and average utility

Marginal Utility= (change in total utility)/(change in quantity consumed)


Average Utility= total utility divided by total quantity consumed.


Quantity Consumed Total Utility Marginal utility Average utility

1 20 ="\frac{20}{1}" =20 "=\frac{20}{1}=20"


2 35 ="\frac{35-20}{2-1}" =15 "=\frac{35}{2}=17.5"


3 47 ="\frac{47-35}{3-2}" =12 "=\frac{47}{3}=15.67"


4 55 ="\frac{55-47}{4-1}" =8 "=\frac{55}{4}=13.75"


5 60 ="\frac{60-55}{5-4}" =5 "=\frac{60}{5}=12"